The newest model of smart car is supposed to get excellent gas mileage. A thorough study showed that gas mileage (measured in miles per gallon) is normally distributed with a mean of 75 miles per gallon and a standard deviation of 10 miles per gallon. What value represents the 50th percentile of this distribution? 85 105 95 75
Answer #1

Given that,

mean = = 75

standard deviation = = 10

Using standard normal table,

P(Z < z) =50 %

= P(Z < z) = 0.50  

= P(Z <0 ) = 0.50

z = 0 Using standard normal table,

Using z-score formula  

x= z * +

x= 0*10+75

x= 75